OpenStudy (anonymous):

This was on the test today and I was so clue else. Ok. It said to put 2y+4x=6 into ax+by=c equation and find A B and C

OpenStudy (anonymous):

IF you put ax+by=c, you already have the answer in front of you. a=4 because it's connected to x b=2 because it's connected to y c=6 because it doesn't have any connection like the ax+by=c equation you were given hope this helps!

OpenStudy (anonymous):

So I don't need to move anything around. It's not in the equation

OpenStudy (anonymous):

Pretty much it shows itself, you can switch 4x and 2y to make it look like the equation if you wanted to

OpenStudy (anonymous):

What if the 4 was negative?

OpenStudy (anonymous):

It's still in the equation, you just put a negative in front of the for, like this: -4x+2y=6 Or 2y-4x=6 or you could write it like this 2y+(-4x)=6

OpenStudy (anonymous):

Oo. Ok thanks (:

OpenStudy (anonymous):

Yup, no problem :)